# Large Language Models as Neurolinguistic Subjects: Identifying Internal Representations for Form and Meaning

### v1

@@ 5.3 Meaning Encoding across Different Forms
How does LLMs’ understanding of meaning change when the form (language) varies? Since our COMPS-ZH and COMPS-DE datasets align with the concepts in the English COMPS dataset, we can explore whether LLMs’ grasp of different linguistic forms for the same concept correlates with their understanding of meaning across languages.
Figure 8: Minimal pair probing results of all three languages.
Our previous results and the detailed results in Figures 15 and 16 suggest that instruction tuning has little influence on the internal representations. Therefore we focus on the base LLMs in the following analysis.
@@
From Figure 8, for all models and languages, form consistently achieves higher F1 scores than meaning, indicating it’s easier for LLMs to make stronger grasp of structural elements compared to conceptual comprehension.
Form needs less data to capture compared to meaning.
From Table 2, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
Figure 9: Meaning competence of LLMs is correlated to their form competence.
@@
Figure 9 illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs.The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
@@ 6 Discussion
Form and Meaning: observations from Saussure’s semiotics Through various experiments focused on Saussure’s concepts of the signifier (form) and the signified (meaning), we have identified several key phenomena. First, LLMs consistently learn linguistic form before they grasp meaning. This suggests a developmental trajectory where statistical patterns in syntax and grammar are more readily captured by the model than conceptual understanding. Second, the models’ formal competence is generally superior to their semantic competence. This is evident in their ability to decode grammaticality structures accurately but with less reliable conceptual accuracy.
Moreover, we observe a linear correlation between form and meaning competence, especially when linguistic forms vary across languages while the underlying meaning remains constant. This pattern implies that an LLM’s understanding of meaning is heavily dependent on its grasp of form. In cases where the form changes (such as between languages), meaning appears to follow the variations in form, indicating that conceptual representation in LLMs is anchored to formal structures rather than an independent comprehension of meaning. These findings highlight the models’ reliance on statistical correlations for understanding meaning, further emphasizing their dependence on formal patterns rather than direct, intrinsic understanding of natural language semantics.
These results help explain the phenomenon of LLMs generating “confidently incorrect” responses, commonly referred to as hallucination, has been a significant issue for a long time Ji et al. (2023). Our results provide a theoretical explanation for this phenomenon from a semiotic and neurolinguistic perspective. Our findings also align with previous studies Mahowald et al. (2024), which suggests that LLMs have good formal competence, but their performance on functional competence (understanding language in the world) is spotty.
Developmental difference between human and machine intelligence. From a perspective of developmental psychology, human kids typically acquire conceptual understanding before mastering grammar Bloom (2002); Tomasello (2005). Pinker (2009)’s semantic bootstrapping hypothesis posits that children initially learn vocabulary through semantic information and then use this semantic knowledge to infer syntactic structures. In contrast, our results indicate that LLMs learn grammar before meaning. Human intelligence is a combination of statistical inference and causal reasoning, whereas LLMs’ intelligence is more likely a result of statistical inference Tenenbaum et al. (2011); Gopnik and Wellman (2012); Lake et al. (2017). Given this nature, the fact that LLMs learn form first might be because grammatical patterns are easier to statistically capture compared to meaning.
Symbol grounding problem and the quest for human-like intelligence
@@ 7 Conclusion
By analyzing Large Language Models (LLMs) as both psycholinguistic and neurolinguistic subjects, we find that LLMs excel at encoding linguistic form but show limitations in understanding meaning. Their grasp of meaning is largely dependent on statistical associations with form. By introducing a cognitive neuroscience perspective, along with semiotics, we hope will inspire further research to deepen our understanding of the language capabilities of LLMs.
@@
We didn’t include experiments covering more languages, which might limit the generalizability of our findings.
The evaluation results for German are poor, which might be attributed to the issue of very long sentences in the DistilLingEval dataset.
Due to limitations in computational resources, we did not perform our experiments on larger-size LLMs, which may lead our results to be biased towards small-scale models. Further studies should include larger models to validate and potentially generalize our findings.
@@
Our project may raise awareness within the computational linguistics community about those low-resource languages.
Appendix A More Psycholinguistic and Neurolinguistic Paradigms’s Cognitive Science Background
@@
Our minimal pair probing is inspired by cognitive neuroscience. In the field of neurolinguistics, decoding analysis has become a fundamental technique in cognitive neuroscience. It tries to extract information encoded in neural patterns (Kriegeskorte et al., 2006). It trains a classifier to predict properties of the stimulus (e.g. a particular image or word input), from the neural responses. If the accuracy of the trained classifier is significantly better than chance, we conclude that the neural data encodes information about the predicted stimulus properties (Norman et al., 2006; Haynes and Rees, 2006).
@@ B.2 Meaning: COMPS, COMPS-ZH, and COMPS-DE
COMPS (Misra et al., 2023) is an English dataset of conceptual minimal pairs for testing an LLM’s knowledge of everyday concepts (e.g., a beaver/*gorilla has a flat tail). This dataset contains 49,340 sentence pairs, constructed using 521 concepts and 3,592 properties. Concepts in the pairs constitute 4 types or knowledge relationship: taxonomy, property norms, co-occurrence and random.
@@
Appendix C More Detailed Minimal Pair Probing Results for English
Figure 10: English minimal pair probing results.
We averaged the results of 4 tasks from the same linguistic level, respectively, and show the result curves in Figure 10. Some noteworthy points from this figure include:
1) The average performances on morphological tasks (purple curve) and on tasks from the semantics-syntax interface (yellow curve) are lower than those for syntactic tasks (blue curve), suggesting that more abstract linguistic levels are easier for LLMs to learn, which also aligns with He et al. (2024). For the ellipsis task especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
@@ Appendix C More Detailed Minimal Pair Probing Results for English
Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much, but will learn concepts better and faster.
Figure 11: Performance difference between Llama3 and Llama2.
Figure 12: Layer-wise t-test results. T-tests were first performed separately on each linguistic task, and then Stouffer’s Z-score method Stouffer et al. (1949) was employed to aggregate the final p-value at the condition level.
@@
Appendix D Multilingual Analysis
@@
4) Poor Performance for German. For German, a low-resource language, all three models perform poorly. Despite Chinese not being a primary training language for the Llama models, their performance is relatively decent, suggesting that the actual proportion of German training data might be much smaller. This highlights differences in the resource allocation for the three languages.
@@
Figure 15: Detailed Chinese decoding results on 6 models. Notice that the pink, orange, blue curves don’t denote morphology or semantics as those in English do. They are made just to make it easier to distinguish in the figure. All non-red curves represent grammatical tasks and red curves represent conceputal tasks.

### v2

@@ 5.3 Multilingual analysis
How does LLMs’ understanding of meaning change when the form (language) varies? Since our COMPS-ZH and COMPS-DE datasets align with the concepts in the English COMPS dataset, we can explore whether LLMs’ grasp of different linguistic forms for the same concept correlates with their understanding of meaning across languages. Our previous results suggest that instruction tuning has little influence on the internal representations. Therefore we focus on the base LLMs here.
@@
From Figure 5-(f), for all models and languages, form consistently achieves higher performance than meaning, indicating it’s easier for LLMs to make a stronger grasp of structural elements compared to conceptual comprehension. Extended multilingual results can be found in Appendix E.
@@
Figure 5-(g) illustrates the relationship between Form competence (x-axis) and Meaning competence (y-axis) across English, German, and Chinese in the neuro assessment for LLMs. The positive correlation (R² = 0.48) suggests that higher meaning competence generally corresponds to higher form competence.
@@ 6 Discussion
Language performance vs. competence: using neurolinguistic probing to reveal deeper competencies.
Our results demonstrate distinct patterns between the neurolinguistic and psycholinguistic evaluation methods. We observed that LLM performance evaluated using the Meta method was significantly lower than that of Direct, while Direct also performed notably worse than the Neuro method. This finding extends Hu and Levy (2023), who compared direct probability measurement (Direct) and meta-linguistic prompting (Meta) and argued that performance does not always reflect competence.
While Hu and Levy (2023) suggested that Direct probability measurement is a more reliable approach for assessing competence, our findings reveal that there is still a significant gap between Direct and Neuro probing, which directly examines internal representations. Hence, we recommend Neuro probing for evaluating competence.
Neuro probing further validates the discrepancy between performance and competence. While Meta results underperform, this does not necessarily indicate that the LLMs lack the underlying linguistic competence. Instead, it may reflect limitations in information access, as suboptimal prompts can prevent models from exhibiting their full capabilities. Specifically, prompting failures do not always equate to a lack of encoded knowledge. This observation is consistent with previous studies Firestone (2020); Lampinen (2024) that emphasize the importance of carefully controlling performance conditions for both humans and machines. methods can often reveal deeper competencies within the models.
Form and meaning: observations from Saussure’s semiotics
Our results reveal that LLMs consistently learn linguistic form before they grasp meaning. This may suggest a developmental trajectory where statistical patterns in syntax and grammar are more readily captured by the model than conceptual understanding. Second, the models’ formal competence is generally superior to their semantic competence. This is evident in their ability to decode grammaticality structures accurately but with less reliable conceptual accuracy.
We further observe a linear correlation between form and meaning competence, particularly when linguistic forms vary across languages while meaning remains constant. This suggests that LLMs’ understanding of meaning might rely heavily on form, with conceptual representation anchored to formal structures rather than independent meaning comprehension. These results offer a semiotic and neurolinguistic explanation for LLMs’ long-standing issue of generating “confidently incorrect” responses, i.e. hallucinations Ji et al. (2023).
@@ 7 Conclusion
This study adopts both psycho- and neuro-linguistic approaches to evaluating LLMs, revealing a distinction between linguistic performance and competence. Our results highlight the limitations of LLMs’ semantic understanding and the need for future research to move beyond statistical correlations toward more grounded language representations. By introducing a cognitive neuroscience perspective, along with semiotics, we hope will inspire further research to deepen our understanding of the language capabilities of LLMs.
@@
This study has several limitations that may impact the generalizability and comprehensiveness of our findings. First, we did not include experiments covering a wider range of languages, which restricts the cross-linguistic applicability of our results. Especially for the analysis and discussion in Section 5.3 on multilingual content, which highlights the necessity of constructing multilingual conceptual datasets.
Second, the evaluation results for German are notably poor, potentially due to the presence of very long sentences in the DistilLingEval dataset, which might have introduced challenges for the models. This underscores the need for constructing syntactic minimal pair datasets for German.
Additionally, our experiments were conducted using small-scale LLMs due to computational resource constraints. This may have introduced a bias in our findings, as larger-scale models could exhibit different behaviors. Future studies should explore larger models to validate and extend the generalizability of these results.
Lastly, the COMPS dataset used for assessing conceptual understanding is not sufficiently fine-grained, as it is limited to only four types of conceptual relationships. A more granular dataset could provide deeper insights into the nuances of how LLMs encode and process meaning. Future work should address this limitation by incorporating more diverse and detailed datasets.
@@
Our project has the potential to raise greater awareness within the computational linguistics community about the challenges faced by low-resource languages. By highlighting the unique linguistic features and limited computational tools available for these languages, we aim to inspire further research and the development of more inclusive language technologies that can better serve underrepresented linguistic communities.
Appendix A Extended Cognitive Science Background
@@
Our minimal pair probing is inspired by cognitive neuroscience. In the field of neurolinguistics, decoding analysis has become a fundamental technique in cognitive neuroscience. It tries to extract information encoded in neural patterns (Kriegeskorte et al., 2006). It trains a classifier to predict the properties of the stimulus (e.g. a particular image or word input), from the neural responses. If the accuracy of the trained classifier is significantly better than chance, we conclude that the neural data encodes information about the predicted stimulus properties (Norman et al., 2006; Haynes and Rees, 2006).
@@ B.2 Meaning: COMPS, COMPS-ZH, and COMPS-DE
COMPS (Misra et al., 2023) is an English dataset of conceptual minimal pairs for testing an LLM’s knowledge of everyday concepts (e.g., a beaver/*gorilla has a flat tail). This dataset contains 49,340 sentence pairs, constructed using 521 concepts and 3,592 properties. Concepts in the pairs constitute 4 types of knowledge relationships: taxonomy, property norms, co-occurrence, and random.
@@
Appendix C Model Details
Resource Level Llama2 Llama3 Qwen
English High High High
Chinese Mid Mid High
German Low Low Low
Llama2 and Llama3
QWen
Appendix D Extended Neurolinguistic Results for English
Some noteworthy points from Figure 5-(a) include:
1) For the ellipsis task especially, local features in layer 0 (the embedding layer before the Transformer structure) already provide sufficient linguistic information to accomplish the task without any contextual information.
@@ Appendix D Extended Neurolinguistic Results for English
Compared to Llama2, Llama3 won’t improve internal grammatical capabilities much but will learn concepts better and faster.
Figure 6: Top: Performance difference between Llama3 and Llama2. Bottom: Layer-wise t-test results. T-tests were first performed separately on each linguistic task, and then Stouffer’s Z-score method Stouffer et al. (1949) was employed to aggregate the final p-value at the condition level.
@@
Appendix E Extended Neurolinguistic Results for Multilingual Analysis
@@
4) Poor Performance for German. For German, a low-resource language, all three models perform poorly. Despite Chinese not being a primary training language for the Llama models, their performance is relatively decent, suggesting that the actual proportion of German training data might be much smaller. This highlights differences in the resource allocation for the three languages. 5) Form needs less data to capture compared to meaning. From Table 3, Chinese is classified as a mid-resourced language for Llama, yet it achieves high form competence (but low meaning competence), suggesting that capturing form requires less data than meaning.
Appendix F Extended Discussion
Developmental difference between human and machine intelligence.
From a perspective of developmental psychology, human kids typically acquire conceptual understanding before mastering grammar Bloom (2002); Tomasello (2005). Pinker (2009)’s semantic bootstrapping hypothesis posits that children initially learn vocabulary through semantic information and then use this semantic knowledge to infer syntactic structures. In contrast, our results indicate that LLMs learn grammar before meaning. Human intelligence is a combination of statistical inference and causal reasoning, whereas LLMs’ intelligence is more likely a result of statistical inference Tenenbaum et al. (2011); Gopnik and Wellman (2012); Lake et al. (2017). Given this nature, the fact that LLMs learn form first might be because grammatical patterns are easier to statistically capture compared to meaning.
Symbol grounding problem and the quest for human-like intelligence
@@ Appendix F Extended Discussion
Figure 9: Detailed Chinese decoding results on 6 models. Notice that the pink, orange, and blue curves don’t denote morphology or semantics as those in English do. They are made just to make it easier to distinguish in the figure. All non-red curves represent grammatical tasks and red curves represent conceptual tasks.